A 79.9g sample of bromine contains approximately 1 mole of bromine atoms. How many bromine atoms are in a .80kg sample of bromine?
1
Expert's answer
2020-09-04T05:18:51-0400

At first, kilogrammes should be turned into grams:

m=80×1000=80000(g).

Then, mass should be divided by molar mass:

n=m/M=80000/79.9=1001.3(mol).

(The data is not very clear. If the point is not mistake, and there are .80 (or 0.80 kg), this value should be used instead:

m=0.80×1000=800(g)

n=800/79.9=10.0(mol)).